### v1

# Tacit Learning with Adaptive Information Selection for Cooperative Multi-Agent Reinforcement Learning

Abstract.
In multi-agent reinforcement learning (MARL), the centralized training with decentralized execution (CTDE) framework has gained widespread adoption due to its strong performance. However, the further development of CTDE faces two key challenges. First, agents struggle to autonomously assess the relevance of input information for cooperative tasks, impairing their decision-making abilities. Second, in communication-limited scenarios with partial observability, agents are unable to access global information, restricting their ability to collaborate effectively from a global perspective. To address these challenges, we introduce a novel cooperative MARL framework based on information selection and tacit learning. In this framework, agents gradually develop implicit coordination during training, enabling them to infer the cooperative behavior of others in a discrete space without communication, relying solely on local information. Moreover, we integrate gating and selection mechanisms, allowing agents to adaptively filter information based on environmental changes, thereby enhancing their decision-making capabilities. Experiments on popular MARL benchmarks show that our framework can be seamlessly integrated with state-of-the-art algorithms, leading to significant performance improvements.
Key words and phrases:
Multi-agent Reinforcement Learning, Adaptive Information Selection, Tacit Learning
1. Introduction
Cooperative Multi-Agent Reinforcement Learning (MARL) has emerged as a robust framework for addressing practical challenges across various domains, including autonomous driving Zhou et al. (2020), gaming Berner et al. (2019), swarm robotics Mguni et al. (2018); Mguni et al. (2019), and smart grids Qiu et al. (2021, 2023); Wang et al. (2021). Despite its success, learning complex cooperative strategies remains a major challenge. Firstly, neglecting the influence of other agents on the system introduces non-stationarity from the perspective of an individual, potentially leading to environmental instability. Additionally, as the number of agents increases, the observation space for joint actions expands exponentially, which may impede the learning process. To effectively address these challenges, the approach of Centralized Training and Decentralized Execution (CTDE) has been proposed and gained popularity in MARL. CTDE utilizes global information during training while achieving decentralized decision-making based on local information. It serves as the foundation for several prominent methods, including VDN Sunehag et al. (2017), MADDPG Lowe et al. (2017), QMIX Rashid et al. (2018), and COMA Foerster et al. (2017).
Despite making progress, CTDE methods still face two major challenges. Firstly, agents in CTDE often rely on specific information for decision-making, and an excessive amount of information may overwhelm them. Secondly, many existing CTDE methods assume mutual independence among agents during training, only incorporating global information in the mixing network. This approach overlooks the importance of information sharing among agents Wang et al. (2022), thus hindering the establishment of cooperation. Therefore, addressing how agents handle and share information in MARL remains a pressing challenge.
Various explicit communication methods have been proposed to address the challenge of information sharing in multi-agent systems, including CommNet Sukhbaatar et al. (2016), BicNet Peng et al. (2017), and NDQ Wang et al. (2019a). To reduce the delays and costs associated with these communication methods, some approaches have shifted toward implicit communication frameworks, which aim to simulate communication by promoting mutual understanding among agents Li et al. (2023); Zhou et al. (2023). However, in these implicit methods, agents often depend too heavily on others to evaluate information, rather than autonomously assessing its importance, effectively delegating decision-making to other agents. A potential solution to this issue is to allow agents to communicate during the training phase while relying solely on local observations during decision-making.
Constrained by cognitive limitations and individual perspectives, humans exhibit selectivity when receiving information. They process this information based on their knowledge and past experiences, selecting the most relevant details for the present moment. In collaborative settings, individuals often develop a tacit understanding through specific training, enabling them to accurately predict and comprehend their peers’ intentions without explicit communication. Inspired by human information processing and cooperation patterns, we propose a novel framework called Selective Implicit Collaboration Algorithm (SICA) for multi-agent systems. SICA is built upon the QMIX framework and can be extended to various methods based on CTDE paradigm. The framework comprises three key blocks: the Selection Block, the Communication Block, and the Regeneration Block. During training, the Selection Block assists agents in filtering information relevant to cooperation, which is then shared with other agents through the Communication Block to generate true information. Subsequently, the Regeneration Block utilizes local information to regenerate true information. Through iterative training, SICA gradually reduces reliance on true information, transitioning from a centralized to a decentralized framework.
Overall, SICA offers several advantages. First, SICA facilitates adaptive selection of input information, ensuring that the obtained information is highly relevant to decision-making, thereby enhancing the performance of CTDE methods. Second, SICA’s exchanges do not require explicit communication, thereby avoiding issues like latency induced by communication and enabling deployment in diverse environments. Finally, SICA serves as a plug-and-play framework for agent policy networks, capable of seamlessly integrating with any value decomposition algorithm.
We evaluated the performance of SICA in the StarCraft Multi-Agent Challenge (SMAC) Samvelyan et al. (2019), SMACv2 Ellis et al. (2022), and Google Research Football (GRF) Kurach et al. (2019). The experimental results demonstrate that SICA outperforms traditional CTDE methods and explicit communication methods in terms of performance.
2. Related Work
CTDE Framework The Centralized Training and Decentralized Execution (CTDE) framework presents a novel MARL approach aimed at addressing the scalability challenges of centralized learning Claus and Boutilier (1998) and the environmental non-stationarity of decentralized learning Tan (1993). This framework operates by training agents based on global information while making decisions based on local information. Two prominent works based on this framework are VDN Sunehag et al. (2017) and QMIX Rashid et al. (2018), which introduce value decomposition methods. VDN aggregates the value functions of individual agents by directly summing them to obtain the joint action function, while QMIX further optimizes this process by ensuring the Individual-Global Maximum (IGM) condition through monotonicity. More recent methods, such as QPLEX Wang et al. (2020), introduce double adversarial networks to overcome representation limitations. Additionally, CTDE encompasses a series of centralized critic methods, including MADDPG Lowe et al. (2017), COMA Foerster et al. (2017), FOP Zhang et al. (2021), etc., which do not impose any restrictions on the representation of the joint action-value function.
Continous Communication CTDE methods typically incorporate global information solely in the mixing network, with agents limited to observing only local information. To augment the decision-making capabilities of agents, numerous studies have adopted communication methods to introduce global information to the agents. One pioneering work in this domain is RIAL with DIAL Foerster et al. (2016), which aims to facilitate communication in CTDE through predefined topological structures. Meanwhile, CommNet was introduced as the first differentiable framework in MARL Sukhbaatar et al. (2016). Subsequent works, such as ATOC Jiang and Lu (2018), I2C Ding et al. (2020)), and DHCG Sukhbaatar et al. (2016), utilize state-dependent communication graphs to address these challenges. ATOC and I2C also introduce gating mechanisms to regulate communication links between agents. Recently, novel communication methods such as TarMAC Das et al. (2018)and NDQ Wang et al. (2019a)have been proposed. TarMAC employs multi-layer attention modules for multi-round communication to strengthen connections between agents. NDQ aims to reduce communication overhead by learning nearly decomposable value functions.
Communication-Free Execution In certain scenarios, factors such as delays and costs can impede the practical deployment of communication methods. To address this challenge, some approaches permit communication only during the training phase to establish cooperative patterns among agents, while prohibiting communication during the execution phase. Previous work introduced a multi-agent framework that utilizes actions as a form of implicit communication, successfully applying it to various robotic tasks Knepper et al. (2017). Building on this concept, researchers developed the PBL framework Tian et al. (2018), which fosters implicit communication through actions and introduces a secondary reward to incentivize this behavior. Furthermore, PBL incorporates a social influence reward specifically for Sequential Social Dilemma (SSD) multi-agent environments Jaques et al. (2018), enabling agents to cooperate more effectively in SSD scenarios.
Our Works While the aforementioned communication-free execution methods rely on gradient flow to facilitate cooperation—thereby increasing model complexity—our approach shares similarities with COLA Xu et al. (2022) and TACO Li et al. (2023). Both methods convey information through specific communication protocols and gradually transition to a decentralized framework. Additionally, to address the challenge of agent information filtering, we designed an adaptive selection mechanism, which effectively enhances the agents’ information processing capabilities.
3. Preliminary
Problem Formulation In a multi-agent cooperative environment, our work adheres to the definition of the Decentralized Partially Observable Markov Decision Process (Dec-POMDP) Bernstein et al. (2000), denoted as ($\mathcal{N}$, $\mathcal{S}$, $U$, $O$, $\Omega$, $P$, $r$, $\gamma$). Here, $\mathcal{N}$ represents the set of $n$ agents, and $\mathcal{S}$ is the global state space of the environment. The joint action space $\boldsymbol{u}=\{u_{1},u_{2},...,u_{n}\}\in\boldsymbol{U}=U^{n}$ consists of the independent actions of each agent, where $\gamma$ represents the discounted factor. In the Partially Observable Markov Decision Process (POMDP), agents lack access to global information. Consequently, at each time step $t$, each agent $i$ can only observe that $o_{i}\in O$ based on the observation function $\Omega(s,u)$: $\mathcal{S}\times\mathcal{U}\rightarrow O$. Agents then select an action $u_{i}$ according to the policy $\pi(u_{i}|o_{i})$. Subsequently, the state transition function $P(s^{\prime}|s,\boldsymbol{u})$ is updated to the next state $s^{\prime}$. Agents collectively receive global rewards according to a reward function $r(s,\boldsymbol{u})$ and share them. The objective for all agents is to find an optimal joint policy $\pi^{*}$ to maximize the global expected return. Our aim is to achieve this goal by designing effective implicit communication between the agents.
Figure 1. A case study in selection mechanism. The colored sections signify the information that the agent elects to remember, whereas the white sections denote the information that the agent chooses to ignore.
The S6 Layer  Given an input scalar $x(t)$, we consider a continuous-time invariant State Space Model (SSM) defined by the following first-order differential equation:
$\dot{h}(t)=Ah(t)_{k-1}+Bx(t)_{k},\quad y(t)=Ch(t)_{k-1}+Dx(t)_{k}$ (1)
Here, $x(t)$ is an input function mapped to produce the output $y(t)$. Previous research has shown that the SSM can capture remote dependencies by initializing matrix $A$ using the HIPPO matrix Gu et al. (2020). Similarly, $D$ is interpreted as a parameter-based skip connection and set to 0, following prior work Gu et al. (2021); Gupta and Berant (2022). Additionally, as the SSM operates on continuous sequences, it is discretized using the Zero-Order Hold method, considering the discrete matrices $\bar{A}$ and $\bar{B}$:
$\bar{A}=\exp(\Delta A),\quad\bar{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B$ (2)
From this, we can rewrite the definitional formula for SSM:
$h_{k}=\bar{A}h_{k-1}+\bar{B}x_{k},\quad y_{k}=Ch_{k}$ (3)
At this point, the parameters $A$, $B$, $C$, $\Delta$ are static, but they are transformed into dynamic parameters to integrate a selection mechanism: $S_{B}(x)=Linear_{N}(x)$, $S_{C}(x)=Linear_{N}(x)$, $S_{\Delta}(x)=Linear_{D}(x)$, $\tau_{\Delta}=softplus$. $Linear$ refers to linear projection. This adjustment enables the model to filter out irrelevant information and retain relevant information over time, enhancing its ability to effectively address the problem.
Value Decomposition  The introduction of value decomposition methods aims to balance the fitting capability of the Q-function with the simplicity of finding its maximum value. Currently, most value decomposition methods adhere to the Individual-Global-Max (IGM) principle Son et al. (2019):
$\begin{split}\mathop{\arg\max}\limits_{\boldsymbol{u}}Q_{joint}(\boldsymbol{\tau},\boldsymbol{u})=\left(\mathop{\arg\max}\limits_{\boldsymbol{u_{1}}}Q_{1}(\tau_{1},u_{1}),\cdots,\mathop{\arg\max}\limits_{\boldsymbol{u_{n}}}Q_{n}(\tau_{n},u_{n})\right)\end{split}$ (4)
This implies that the joint action, which is composed of local greedy actions chosen by each agent based on their individual Q-functions, is equivalent to the joint greedy action chosen based on $Q_{joint}$. Two classic value decomposition baselines are VDN Sunehag et al. (2017) and QMIX Rashid et al. (2018). Taking QMIX as an example, the estimated value $Q_{tot}$ adopts the design of a monotonic mixing network design, while each agent has its own policy network. Both networks contain learnable parameters $\theta=\{\theta_{u},\theta_{v}\}$, which can be updated by minimizing the Temporal Difference (TD) Loss:
$\mathcal{L}_{TD}(\theta)=E_{\mathcal{D}}[(y_{tot}-Q_{tot}(\boldsymbol{\tau},\boldsymbol{u},s;\theta))^{2}]$ (5)
Here, $y_{tot}=r+\gamma\max\limits_{\boldsymbol{u^{\prime}}}Q^{\prime}_{tot}(\boldsymbol{\tau}^{\prime},\boldsymbol{u}^{\prime},s^{\prime};\theta^{-}))$, $E[\cdot]$ denotes the expectation function, and $\mathcal{D}$ represents the replay buffer of the transition.
4. SICA
In this section, we delve into the core details of SICA, outlining its adaptive information selection mechanism, communication strategy, and the transition from explicit to implicit communication. We then elucidate the overarching objectives of the agents and the learning process. Notably, our proposed method operates within the agent network and can be seamlessly integrated with any value decompostition method.
4.1. Adaptive information selection
In cooperative multi-agent tasks, a crucial aspect is the agents’ capacity to autonomously filter input information. This involves selecting the most relevant information for ongoing cooperation, guided by experience and task requisites, while minimizing attention to irrelevant data. Currently, there is no CTDE method that equips agents with such discernment. To address this, we propose an adaptive information selection mechanism for agents. This mechanism operates by integrating historical information at each time step, giving more detail to data closer to the current time step while abstracting information that is further away. Subsequently, based on the prevailing cooperation demands, agents discern effective information and filter out irrelevant data. For instance, Figure 1 illustrates this selection mechanism in action within the SMAC environment, where two Marines and one Medivac from the red team engage a blue Medivac. Considering the red Medivac, given the proximity of both factions, it disregards distance information. Additionally, with no units generating shields on the field, the red Medivac also discounts shield information.
4.2. SICA Framework
Figure 2. The overall framework of SICA. We illustrate the network architecture using the example of the $Nth$ agent. The overall architecture comprises a mixing network and agent networks ($left$). Details of the agent network include the Selection Block, Communication Block and Regeneration Block ($middle$). The Selection Block ($right$)
To address the mentioned challenges and facilitate framework extension, SICA adopts the classic QMIX Rashid et al. (2018) framework. This framework comprises a mixing network and agent networks. However, unlike QMIX, SICA does not presume agent independence during training. SICA primarily enhances the agent network, as depicted in the overall framework in Figure 2 left. This framework encompasses the Selection Block, Communication Block, and Regeneration Block, as delineated in Figure 2 middle. The Selection Block empowers the framework with information selection capabilities. Agents can dynamically choose desired input information, as elucidated in Figure 2 right.
Selection Block  The Selection Block consists of two MLPs and an S6 layer Gu and Dao (2023). Since the time intervals of the inputs in the selective tasks are variable, a time-varying model is required, so we integrated the S6 layer into the framework. The two MLPs and other components can be modeled as a Gating Unit (GU) Liu et al. (2021), which is responsible for learning long-term dependencies in the input. Therefore, the entire module can be seen as a dual selection mechanism that combines the gating mechanism with the S6 selection mechanism. Additionally, to empower the Selection Block to thoroughly select information, we establish a mini-buffer preceding it. This mini-buffer preserves the preceding $b$ observation-action pairs of a single agent, constituting a mini-batch $\{(o_{i}^{t-1},u_{i}^{t-2}),\cdots,\\ (o_{i}^{t-b},u_{i}^{t-b-1})\}$ transmitted to the Selection Block. Here, $b$ represents the capacity of the mini-buffer and $i$ represents the agent $i$, both acting as hyperparameters.
Agent $i$’s observation-action pairs from the current time step and the previous $b$ time steps are integrated and passed through a GU before being fed to the S6 layer. Within the S6 layer, $\Delta$, $B$, and $C$ function as dependencies on the input, rendering the parameters data-dependent in SICA. $\Delta$ aids in controlling whether the agent prioritizes the current hidden state or the input, while $B$ and $C$ facilitate the filtering out of irrelevant information. As only the hidden state information from S6 is required, solely $B$ is utilized. The computation process for the entire module is outlined as follows:
$z_{i}^{t}$ $=\text{MLP}(x_{i1}^{t})\sigma(\text{MLP}(x_{i2}^{t}))$
$h_{i}^{t}$ $=\bar{A}h_{i}^{t-1}+\bar{B}z_{i}^{t}$
Here, $\{x_{i1}^{t},x_{i2}^{t}\}=\text{Split}(x_{i}^{t})$. The discrete matrices $\bar{A}$ and $\bar{B}$ can be derived from Equation 2.
Communication Block The Communication Block complements the Selection Block by enabling agents to incorporate global information into their decision-making process. In this block, agents prioritize suggestions from other agents and receive information through attention-weighted mechanisms Vaswani et al. (2017). Given the hidden states of agents $i$ and $j$ as input, we consider two learnable matrices: the self-query matrix $q_{i}^{t}=W_{q}h_{i}^{t}$ and the cognition matrix $k_{j}^{t}=W_{k}h_{j}^{t}$, where $W_{q}$ , $W_{k}$ are both learnable linear transformations. The calculation of attention weights proceeds as follows:
$c_{i,j}^{t}$ $=\frac{(q_{i}^{t})^{T}k_{j}^{t}}{\sqrt{d_{h}}}$ (7)
$w_{i,j}^{t}$ $=\frac{\exp(c_{i,j}^{t})}{\sum_{k=1}^{N}\exp(c_{i,k}^{t})}$ (8)
$d_{h}$ represents the dimension of the hidden state, and $\exp$ denotes the exponential operation. Then the true information $v_{i}^{t}$ obtained by agent $i$ can be calculated as $v_{i}^{t}=\sum_{i\neq j}w_{i,j}h_{j}^{t}$.
Figure 3. Performance comparison between SICA and baselines on SMAC.
Regeneration Block To ensure that decision-making relies solely on local information, we must convert the existing centralized framework into a decentralized one. Hence, we introduce the Regeneration Block, comprising the Selection Block and an MLP, which exclusively takes the agent $i$’s observation-action pair as input, as depicted in the central section of Figure 2. This Regeneration Block allows us to derive the regenerated information $\hat{v}_{i}$, which continuously approximates the true information $v_{i}$. It’s notable that, during the initial stages of training, to ensure accurate learning of global information by the Regeneration Block, we include observation-action pairs from other agents in its minibuffer, gradually reducing it to 00 over time.
We utilize exponential weighted averaging to ensure that the regenerated information $\hat{v}_{i}^{t}$ converges towards the true information $v_{i}^{t}$. Then, we compute the cross-information $\bar{v}_{i}^{t}$ by weighting both of them.
$\bar{v}_{i}^{t}=(1-\alpha(t))\hat{v}^{t}+\alpha(t)v_{i}^{t}$ (9)
Here, $\alpha(t)$ is dynamic, and to ensure a smoother transition of the framework, we update it using a method similar to cosine annealing:
$\alpha(t)=\alpha_{start}+(\alpha_{final}-\alpha_{start})\cos(\frac{t}{t_{\max}}\pi).$ (10)
We set $\alpha_{start}$ to 1 and $\alpha_{final}$ to 0.
As the training progresses, the cross information $\bar{v}_{i}^{t}$ gradually transitions into regenerated information, allowing the framework to rely less on true information after training completion. Finally, we concatenate the $\bar{v}_{i}^{t}$ with $h_{i}^{t}$ and pass it through an MLP block, obtaining the action value $Q_{i}$ of agent $i$.
Through the Selection Block and Communication Block, agents effectively utilize information, while the Regeneration Block facilitates the transition from centralized to decentralized decision-making. Overall, compared to other CTDE methods, SICA achieves a balance between decision-making capability and generality.
4.3. Learning Objective
The overall learning objective of our method is divided into two parts: the TD loss function, which constitutes the end-to-end optimal value decomposition, and the minimization of the regeneration information error.
The TD loss function part is the same as in QMIX in Equation 5 and has the following form:
$\mathcal{L}_{TD}(\theta)=E_{\mathcal{D}}[(r+\gamma\max\limits_{\boldsymbol{u}^{\prime}}Q^{{}^{\prime}}_{tot}(\boldsymbol{\tau}^{\prime},\boldsymbol{u}^{\prime},s^{\prime};\theta^{-})-Q_{tot}(\boldsymbol{\tau},\boldsymbol{u},s;\theta))^{2}]$ (11)
Here, $\theta$ represents the learnable parameters, $\theta^{-}$ are the parameters of a target network as in DQN, $E[\cdot]$ denotes the expectation function, and $\mathcal{D}$ represents the replay buffer of transitions.
The part of minimizing the regenerated information error is to ensure that the regenerated information $\hat{v}_{i}$ closely approximates the true information $v_{i}$, facilitating the smooth transition of SICA from a centralized architecture to a decentralized one. To achieve this objective, we introduce an auxiliary loss function called alignment loss function $\mathcal{L}_{Align}$ , formulated as follows:
$\mathcal{L}_{Align}(\theta)=\frac{1}{n}\sum\limits_{i=1}^{n}E[(\hat{v}_{i}(h_{i};\theta^{-})-v_{i}(\boldsymbol{h};\theta))^{2}]$ (12)
Where $E[\cdot]$ denotes the expectation function and $n$ denotes the number of agents.
These two parts are trained simultaneously, but with a gradual increase in the importance of the alignment loss function $\mathcal{L}_{Align}$ to ensure the performance of the architecture. The total loss function can be expressed as follows:
$\mathcal{L}_{tot}(\boldsymbol{\tau},\boldsymbol{u},s,h_{i},\boldsymbol{h};\theta)=\mathcal{L}_{TD}+\sigma(t)\mathcal{L}_{Align}$ (13)
$t$ denotes a time step, and $\sigma(t)$ is a threshold function defined as follows:
$\sigma(t)=\left\{\begin{aligned} \beta_{1}\quad t\leq T\\ \beta_{2}\quad t>T\\ \end{aligned}\right.$ (14)
$T$, $\beta_{1}$ and $\beta_{2}$ are hyperparameters, and they are fixed values that need to be satisfied, i.e., the proportion of reconstruction loss increases as the information increases.
Figure 4. Performance comparison between SICA and baselines on SMACv2.
5. Experiments
[…]
5.1. Performance on Multi-Agent Benchmarks
In this subsection, our goal is to address questions 1 and 2 by assessing SICA’s performance across widely-used MARL benchmarks.
SMAC We initiate our evaluation by assessing SICA’s performance in the SMAC, where our aim is to control a team of allied units against an enemy team governed by built-in policies. Victory is achieved by eliminating all enemy units within a chapter’s time limit. Our metric for evaluation is the alliance team’s win rate, which we aim to maximize. We compare SICA against several robust baselines grounded in the CTDE framework. These include clasSICAl methods like VDN, QMIX, and QTRAN Son et al. (2019), as well as communication-based approaches like QMIX-Attention Hu et al. (2021) and NDQ Wang et al. (2019b). Given that most methods perform adequately in easy maps, we focus on evaluating their performance in more challenging environments, particularly two hard maps ($5m\_vs\_6m$, $2c\_vs\_64zg$) and four super hard maps ($3s5z\_vs\_3s6z$, $MMM2$, $27m\_\\ vs\_30m$, $Corridor$) to offer a comprehensive assessment of SICA’s capabilities.
The median win rates across different maps are depicted in Figure 3. SICA consistently outperforms the baselines across all maps, even surpassing explicit communication methods. This underscores the effectiveness of SICA in information processing and highlights the robustness of its information regeneration capability. Across all methods, there is a noticeable decline in win rates as we transition from hard to super hard maps, which aligns with expectations given the heightened complexity of the latter scenarios. It’s worth mentioning that QTRAN and NDQ exhibit suboptimal performance. QTRAN exhibits suboptimal performance across all maps, potentially attributable to challenges in credit assignment resulting in the development of passive agents. Meanwhile, NDQ demonstrates efficacy solely on select maps, potentially stemming from instability in its message passing methodology.
SMACv2 Next, we evaluate SICA on SMACv2. SMACv2 is an enhanced version of SMAC. In SMACv2, each agent is equipped with randomly generated unit types and initial positions, introducing unpredictability into the environment. These unit types are generated according to a fixed probability distribution, with agents unaware of their own unit types, necessitating adaptable strategies capable of addressing all potential unit types. Figure 4 presents a comparative analysis between SICA and baseline methods on three randomly generated maps for the $Protoss$, $Terran$, and $Zerg$ races. Remarkably, SICA consistently outperformed the baselines, even in highly stochastic SMACv2 environments. This finding underscores the effectiveness of SICA in task completion, highlighting the robustness of its selection mechanism and information regeneration capabilities.
GRF Finally, we evaluated SICA’s performance on GRF, a MARL benchmark based on the open-source game Gameplay Football. In this environment, agents collaborate to orchestrate attacks, with rewards solely granted upon goal scoring. In Table 1, we selected the explicit communication method QMIX-Attention and the current state-of-the-art method CDS-QMIX Li et al. (2021) as baselines. We then compared SICA with these baselines across the two most challenging scenarios. The agents were trained for 10 million steps using 8 threads in all scenarios. The results demonstrate that SICA significantly outperformed the other methods, thereby highlighting its efficacy in diverse environments.
Table 1. Performance comparison between SICA and baselines on GRF
Algorithm $3\_vs\_1\_with\_keeper$ $counterattack\_hard$
CDS-QMIX 0.760 $\pm$ 0.141 0.585 $\pm$ 0.072
QMIX-Attention (Comm) 0.605 $\pm$ 0.182 0.397 $\pm$ 0.114
SICA (Ours) 0.854 $\pm$ 0.061 0.671 $\pm$ 0.141
Figure 5. Learning curves with different numbers of agents in SMACv2 $protoss$.
5.2. Ablation Studies
In this subsection, we will conduct ablation studies. Three sets of experiments were conducted to address questions 3, 4, and 5, respectively. All experiments were performed in the challenging SMACv2 scenario.
Different numbers of agents Given SICA’s ability to encapsulate historical information, we hypothesized that it could effectively handle larger volumes of data, prompting question 3. To answer this question, we tested scenarios with different numbers of agents. We compared SICA with the baseline, QMIX-Attention, the best-performing explicit communication method identified in earlier experiments. As illustrated in Figure 5, SICA consistently outperformed the baseline across all scenarios, highlighting its effectiveness in handling larger-scale multi-agent environments.
Replace Selection Block To address question 4, we introduce a variant of SICA called ICA. In this variant, the Selection is replaced by a gated structure consisting of a MLP Block and a GRU Cell. The results of SICA and ICA in $protoss\_5\_vs\_5$ and $terran\_5\_vs\_5$ scenarios are shown in Figure 6. We observe that the performance of SICA markedly surpassed that of ICA. This suggests that the Selection Block consistently enhances the agents’ ability and underscores the significance of information filtering in multi-agent tasks.
Figure 6. Performance comparison between SICA and ICA.
Fixed $\alpha$ Does SICA really need progressive information regeneration? To answer question 5, we set up two scenarios: one where $\alpha(t)$ is fixed at 0, disregarding $\mathcal{L}_{align}$ , which we call SICA-ZERO; and another where $\alpha(t)$ is fixed at 1 and suddenly switches to 0 near the end of the training, which we call SICA-ONE. For SICA-ZERO, the communication module does not participate, making SICA equivalent to a traditional CTDE method, relying on local information for training. For SICA-ONE, the training process fully depends on communication, forcing the agents to learn complex information directly within a short time. The comparison between these two scenarios and SICA is shown in Figure 7. Regardless of the value of $\alpha(t)$, SICA consistently performs better. Therefore, progressive information regeneration is necessary as it guides the learning process and enhances the agents’ capabilities.
Figure 7. Performance comparison between SICA and SICA’s Variants.
5.3. More Studies
Applying SICA to VDN  To demonstrate the versatility of SICA as a plug-and-play framework, we apply SICA to VDN, referred to as SICA-VDN, in this section. The comparison curve between SICA-VDN and VDN is illustrated in Figure 8, showing that SICA markedly improves the performance of the baseline VDN.
Figure 8. Performance comparison between SICA-VDN and VDN.
Comparing SICA with another communication-free execution method.  In this section, we compare SICA with another communication-free execution baseline, QMIX-CADP Zhou et al. (2023),on SMACv2. QMIX-CADP adopts direct pruning to transition the framework into a decentralized framework. In Figure 9, we can observe that SICA’s performance is significantly better than that of QMIX-CADP, indicating that a gradual framework transition is more reasonable than a direct transition.
Figure 9. Performance comparison between SICA and QMIX-CADP.
(a) Protoss_3_vs_5
(b) Terran_3_vs_5
Figure 10. Visualizations of the regenerated information.
Visualization  To verify whether the Regeneration Block can truly approximate the true information, we visualize the true information and the regenerated information from the $Protoss$ and $Terran$ maps using t-SNE Van der Maaten and Hinton (2008) compression into two-dimensional embeddings. In this section, we adopt the $53\_vs\_$ 5 mode with higher difficulty and randomness to enhance the difficulty of regeneration. As shown in Figure 10, although some bias is present in both scenarios, the regenerated information can reflect the distribution of the true information effectively. Therefore, the regeneration process is effective.
Scalability  We combined SICA with the actor-critic method MADDPG and compared it to the traditional MADDPG algorithm in the $Predator-Prey$, $Navigation$, and $Pantomime$ environments. Table 2 presents the experimental results after 2 million training steps, demonstrating that SICA continues to enhance the overall performance of the framework.
Table 2. Performance comparison between SICA-MADDPG and MADDPG
Algorithm $Predator-Prey$ $Navigation$ $Pantomime$
MADDPG 831.48 $\pm$ 98.73 -233.31 $\pm$ 32.10 -812.79 $\pm$ 592.01
SICA-MADDPG 725.11 $\pm$ 312.26 -255.23 $\pm$ 61.42 -926.55 $\pm$ 601.92
6. Limitations
In this section, we examine three potential limitations of SICA. First, the Regeneration Block may encounter difficulties when there is no correlation between the trajectories or observations of the agents, as it relies on similar historical information for regeneration. Although we mitigate this limitation by incorporating observations from other agents into a mini buffer, there may still be instances where those agents fail to utilize this portion of historical data. Second, while we are confident in SICA’s performance in large-scale multi-agent environments, our experiments are constrained by computational resources, necessitating further evaluation of its effectiveness in larger-scale scenarios. Thirdly, when applying SICA to new tasks, adjustments to multiple hyperparameters will be necessary. Finally, we implemented parameter sharing among the agents to accelerate the training process, as not using this technique would significantly increase the training time.
7. Conclusion
In this paper, we introduced a novel MARL architecture named SICA, designed to enhance agents’ information handling capabilities and improve the framework’s generality. By integrating information selection with communication mechanisms, SICA empowers agents to autonomously choose relevant information while incorporating information from other agents. To accommodate to communication-limited environments, SICA gradually learns the tacit understanding between agents, eventually transitioning to a fully decentralized framework. Experimental results illustrate SICA’s effectiveness in regenerating global information and significantly enhancing performance in challenging multi-agent tasks through information selection. In future work, we aim to seamlessly extend the framework to encompass various CTDE methods.